callc1111:

Kira drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 10 hours. When Kira drove home, there was no traffic and the trip only took 7 hours. If her average rate was 18 miles per hour faster on the trip home, how far away does Kira live from the mountains?

Vocaloid:

average rate = distance/time so trip there: r1 = x/10 so trip back r2 = x/7 so r2 is greater than r1 by 18 mph translate this into an equation and solve for x

Vocaloid:

Closed due to answered.